OpenStudy (anonymous):

ALMOST DONE!!! PLEASE HELP!!! A surveyor locates points A and B at the same elevation and 3950ft apart. At A, the angle of elevation to the summit of the mountain is 18 degrees. At B, the angle of elevation is 31 degrees. a).Find the height of the mountain. (round to the nearest hundredth)

OpenStudy (anonymous):



OpenStudy (anonymous):

tan18=x/3950

OpenStudy (anonymous):

x=3950(tan18) x=1283.43

OpenStudy (anonymous):

how can i find angle C

OpenStudy (anonymous):

like the other side of angle C and B

OpenStudy (anonymous):

Angle B = 180-31

OpenStudy (anonymous):

why is it 180? -31

OpenStudy (anonymous):

angle c = 180-(59+90+18)

OpenStudy (anonymous):

the measurement of horizontal line is always 180 degrees

OpenStudy (anonymous):

angle C= 72 degrees

OpenStudy (anonymous):

but 180-1657=13

OpenStudy (anonymous):

i mean 180-167=13

OpenStudy (anonymous):